# Efficient designs for multivariate crossover trials

### v1

@@ Appendix B Efficiency under the generalized markov-type structure
[…]
$\begin{bmatrix}\bm{Y}^{{}^{\prime}}_{d1}&\ldots&\bm{Y}^{{}^{\prime}}_{dg}\end{bmatrix}^{{}^{\prime}}$ $=\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)\bm{\eta}^{(1)}+\begin{bmatrix}\bm{I}_{g}\otimes\bm{{1}}_{np}&\bm{Z}_{1}&\bm{I}_{g}\otimes\bm{F}_{d}\end{bmatrix}\bm{\eta}^{(2)}+\begin{bmatrix}\bm{\epsilon}^{{}^{\prime}}_{1}&\ldots&\bm{\epsilon}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}}.$ (S.5)
Premultiplying (S.5) by $\bm{\varSigma}^{-1/2}$, we get the model as
$\begin{bmatrix}\bm{Y}^{{}^{\prime}}_{d1(new)}&\ldots&\bm{Y}^{{}^{\prime}}_{dg(new)}\end{bmatrix}^{{}^{\prime}}=\bm{\varSigma}^{-1/2}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)\bm{\eta}^{(1)}+\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{I}_{g}\otimes\bm{{1}}_{np}&\bm{Z}_{1}&\bm{I}_{g}\otimes\bm{F}_{d}\end{bmatrix}\bm{\eta}^{(2)}+\\ \begin{bmatrix}\bm{\epsilon}^{{}^{\prime}}_{1(new)}&\ldots&\bm{\epsilon}^{{}^{\prime}}_{g(new)}\end{bmatrix}^{{}^{\prime}},$ (S.6)
where $\bm{Y}_{dk(new)}=\bm{\varSigma}^{-1/2}\bm{Y}_{dk}$ and $\bm{\epsilon}_{k(new)}=\bm{\varSigma}^{-1/2}\bm{\epsilon}_{k}$. Thus using the expression of the information matrix from Kunert, (1983), we get that under the proportional structure, the information matrix for the direct effects can be expressed as
$\bm{C}_{d(s1)}$ $=\left(\bm{I}_{g}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{Z}_{1}&\bm{I}_{g}\otimes\bm{F}_{d}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right).$ (S.8)
From (S.8) and using the result $pr^{\perp}\left(\begin{bmatrix}\bm{\zeta}_{1}&\bm{\zeta}_{2}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\zeta}_{1}\right)-pr^{\perp}\left(\bm{\zeta}_{1}\right)\bm{\zeta}_{2}\left(\bm{\zeta}^{{}^{\prime}}_{2}pr^{\perp}\left(\bm{\zeta}_{1}\right)\bm{\zeta}_{2}\right)^{-}\times$
$\bm{\zeta}^{{}^{\prime}}_{2}pr^{\perp}\left(\bm{\zeta}_{1}\right)$ (Kunert, 1983), we get
@@
where $\bm{C}_{d(s1)(11)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)$, $\bm{C}_{d(s1)(12)}=\bm{C}^{{}^{\prime}}_{d(s1)(21)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$ and $\bm{C}_{d(s1)(22)}=\left(\bm{I}_{g}\otimes\bm{F}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$. Here, $\bm{A}^{*}=\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\bm{Z}_{1}\right)\bm{\varSigma}^{-1/2}$. ∎
@@
$\bm{A}^{*}$ $=\bm{\varGamma}^{-1}\otimes\left(\bm{H}_{n}\otimes\bm{V}^{*}\right),$ (S.10)
where $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$, and $\bm{\varGamma}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal matrix. Here, $\bm{H}_{n}=\bm{I}_{n}-\frac{1}{n}\bm{{1}}_{n}\bm{{1}}_{n}^{{}^{\prime}}$.
@@
$\bm{\varSigma}$ $=\bm{\varGamma}\otimes\left(\bm{I}_{n}\otimes\bm{V}\right),$ (S.11)
where $\bm{V}$ is a known positive definite and symmetric matrix. $\bm{\varGamma}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal elements.
@@
$\begin{split}\bm{\varSigma}^{-1}&=\bm{\varGamma}^{-1}\otimes\left(\bm{I}_{n}\otimes\bm{V}^{-1}\right),\\ \bm{\varSigma}^{-1}\bm{Z}_{1}&=\bm{\varGamma}^{-1}\otimes\left(\bm{I}_{n}\otimes\bm{V}^{-1}\right)\bm{X}_{1},\text{ and}\\ \bm{Z}^{{}^{\prime}}_{{1}}\bm{\varSigma}^{-1}\bm{Z}_{1}&=\bm{\varGamma}^{-1}\otimes\bm{X}^{{}^{\prime}}_{{1}}\left(\bm{I}_{n}\otimes\bm{V}^{-1}\right)\bm{X}_{1}.\end{split}$ (S.12)
From Bose and Dey, (2009), we get that $\left[\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)pr^{\perp}\left(\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\bm{X}_{1}\right)\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\right]=\bm{H}_{n}\otimes\bm{V}^{*}$, where $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$. Here, $\bm{J}_{p\times p}=\bm{{1}}_{p}\bm{{1}}_{p}^{{}^{\prime}}$ and $\bm{H}_{n}=\bm{I}_{n}-\frac{1}{n}\bm{{1}}_{n}\bm{{1}}_{n}^{{}^{\prime}}$. Thus from (S.12) and using the expression $\bm{A}^{*}=\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\bm{Z}_{1}\right)\bm{\varSigma}^{-1/2}$ as given in Lemma S.1, we get that for $g>1$
$\bm{A}^{*}$ $=\bm{\varGamma}^{-1}\otimes\left[\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)pr^{\perp}\left(\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\bm{X}_{1}\right)\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\right]$
$=\bm{\varGamma}^{-1}\otimes\left(\bm{H}_{n}\otimes\bm{V}^{*}\right).$
@@
$=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{Z}_{1}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$ (S.15)
@@
$\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$ and $\bm{C}_{{d(s2)}(22)(1)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\bm{H}_{t}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$. Here, $\bm{A}^{*}=\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\bm{Z}_{1}\right)\bm{\varSigma}^{-1/2}$ and $\bm{H}_{t}=\bm{I}_{t}-\frac{1}{t}\bm{{1}}_{t}\bm{{1}}_{t}^{{}^{\prime}}$.
@@
$\bm{C}_{d(s2)}$ $=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\begin{bmatrix}\bm{Z}_{1}&\bm{I}_{2}\otimes\bm{F}_{d}\end{bmatrix}\right)\bm{\varSigma}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right).$ (S.18)
From (S.18) and using the result $pr^{\perp}\left(\begin{bmatrix}\bm{\zeta}_{1}&\bm{\zeta}_{2}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\zeta}_{1}\right)-pr^{\perp}\left(\bm{\zeta}_{1}\right)\bm{\zeta}_{2}\left(\bm{\zeta}^{{}^{\prime}}_{2}pr^{\perp}\left(\bm{\zeta}_{1}\right)\bm{\zeta}_{2}\right)^{-}\times$
$\bm{\zeta}^{{}^{\prime}}_{2}pr^{\perp}\left(\bm{\zeta}_{1}\right)$ (Kunert, 1983), we get
@@
where $\bm{C}_{d(s2)(11)}=\left(\bm{I}_{2}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$, $\bm{C}_{d(s2)(12)}=\bm{C}^{{}^{\prime}}_{d(s2)(21)}=\left(\bm{I}_{2}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\right)$ and $\bm{C}_{d(s2)(22)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\right)$. Here, $\bm{A}^{*}=\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\bm{Z}_{1}\right)\bm{\varSigma}^{-1/2}$.
Since $\bm{F}_{d}\bm{{1}}_{t}=\bm{{1}}_{n}\otimes\begin{bmatrix}0&\bm{{1}}_{p-1}^{{}^{\prime}}\end{bmatrix}^{{}^{\prime}}=\bm{P}\begin{bmatrix}0&\bm{{1}}_{p-1}^{{}^{\prime}}\end{bmatrix}^{{}^{\prime}}$, we get $\left(\bm{I}_{2}\otimes\frac{1}{t}\bm{F}_{d}\bm{{1}}_{t}\bm{{1}}_{t}^{{}^{\prime}}\right)=\left(\bm{I}_{2}\otimes\bm{P}\right)\times$
$\left(\bm{I}_{2}\otimes\frac{1}{t}\begin{bmatrix}0&\bm{{1}}_{p-1}^{{}^{\prime}}\end{bmatrix}^{{}^{\prime}}\bm{{1}}_{t}^{{}^{\prime}}\right)$. Hence, $\bm{I}_{2}\otimes\frac{1}{t}\bm{F}_{d}\bm{{1}}_{t}\bm{{1}}_{t}^{{}^{\prime}}$ belongs to the column space of $\bm{I}_{2}\otimes\bm{P}$. So,
$\bm{\varSigma}^{-1/2}\left(\bm{I}_{2}\otimes\frac{1}{t}\bm{F}_{d}\bm{{1}}_{t}\bm{{1}}_{t}^{{}^{\prime}}\right)$ belongs to the column space of $\bm{\varSigma}^{-1/2}\bm{Z}_{1}$. Thus from (S.18), we can prove (S.15).
Further using the result $pr^{\perp}\left(\begin{bmatrix}\bm{\zeta}_{1}&\bm{\zeta}_{2}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\zeta}_{1}\right)-pr^{\perp}\left(\bm{\zeta}_{1}\right)\bm{\zeta}_{2}\left(\bm{\zeta}^{{}^{\prime}}_{2}pr^{\perp}\left(\bm{\zeta}_{1}\right)\bm{\zeta}_{2}\right)^{-}\bm{\zeta}^{{}^{\prime}}_{2}pr^{\perp}\left(\bm{\zeta}_{1}\right)$ (Kunert, 1983), we get
@@
where $\bm{C}_{{d(s2)}(11)(1)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$, $\bm{C}_{{d(s2)}(12)(1)}=\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=$
$\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$ and $\bm{C}_{{d(s2)}(22)(1)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\bm{H}_{t}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$. Here, $\bm{H}_{t}=\bm{I}_{t}-\frac{1}{t}\bm{{1}}_{t}\bm{{1}}_{t}^{{}^{\prime}}$. ∎
@@
$\bm{A}^{*}$ $=\begin{bmatrix}\bm{H}_{n}\otimes\bm{\varOmega}_{1}&-\bm{H}_{n}\otimes\bm{\varOmega}_{2}\\ -\bm{H}_{n}\otimes\bm{\varOmega}_{2}&\bm{H}_{n}\otimes\bm{\varOmega}_{4}\end{bmatrix},$ (S.23)

### v2

@@ Appendix B Efficiency under the generalized Markov-type structure
Let $\bm{\eta}^{(1)}=\begin{bmatrix}\bm{\tau}^{{}^{\prime}}_{1}&\cdots&\bm{\tau}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}}$ and $\bm{\eta}^{(2)}=\begin{bmatrix}\mu_{1}&\mu_{2}&\bm{\alpha}^{{}^{\prime}}_{1}&\bm{\beta}^{{}^{\prime}}_{1}&\cdots&\bm{\alpha}^{{}^{\prime}}_{g}&\bm{\beta}^{{}^{\prime}}_{g}&\bm{\rho}^{{}^{\prime}}_{1}&\cdots&\bm{\rho}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}}$. So $\bm{\eta}^{(1)}$ denotes the vector of direct effects. Then by rearranging the parameters, the model (2.2) can be equivalently expressed as
$\begin{bmatrix}\bm{Y}^{{}^{\prime}}_{d1}&\cdots&\bm{Y}^{{}^{\prime}}_{dg}\end{bmatrix}^{{}^{\prime}}$ $=\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)\bm{\eta}^{(1)}+\begin{bmatrix}\bm{I}_{g}\otimes\bm{{1}}_{np}&\bm{Z_{1}}&\bm{I}_{g}\otimes\bm{F}_{d}\end{bmatrix}\bm{\eta}^{(2)}+\begin{bmatrix}\bm{\epsilon}^{{}^{\prime}}_{1}&\cdots&\bm{\epsilon}^{{}^{\prime}}_{g}\end{bmatrix}^{{}^{\prime}}.$ (S.2)
Premultiplying (S.2) by $\bm{\mathit{\Sigma}}^{-1/2}$, we get the model as
$\begin{bmatrix}\bm{Y}^{{}^{\prime}}_{d1(new)}&\cdots&\bm{Y}^{{}^{\prime}}_{dg(new)}\end{bmatrix}^{{}^{\prime}}=\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)\bm{\eta}^{(1)}+\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{I}_{g}\otimes\bm{{1}}_{np}&\bm{Z_{1}}&\bm{I}_{g}\otimes\bm{F}_{d}\end{bmatrix}\bm{\eta}^{(2)}+\\ \begin{bmatrix}\bm{\epsilon}^{{}^{\prime}}_{1(new)}&\cdots&\bm{\epsilon}^{{}^{\prime}}_{g(new)}\end{bmatrix}^{{}^{\prime}}.$ (S.3)
Thus using the expression of the information matrix from Kunert, (1983), we get that under the proportional structure, the information matrix for the direct effects can be expressed as
$\bm{C}_{d(s1)}$ $=\left(\bm{I}_{g}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{Z_{1}}&\bm{I}_{g}\otimes\bm{F}_{d}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right).$ (S.4)
From (S.4) and using the result $pr^{\perp}\left(\begin{bmatrix}\bm{\mathit{\zeta}_{1}}&\bm{\mathit{\zeta}_{2}}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)-pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)\bm{\mathit{\zeta}_{2}}\left(\bm{\mathit{\zeta}}^{{}^{\prime}}_{\mathbf{2}}pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)\bm{\mathit{\zeta}_{2}}\right)^{-}\times$
$\bm{\mathit{\zeta}}^{{}^{\prime}}_{\mathbf{2}}pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)$ (Kunert, 1983), we get
@@
where $\bm{C}_{d(s1)(11)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)$, $\bm{C}_{d(s1)(12)}=\bm{C}^{{}^{\prime}}_{d(s1)(21)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$ and $\bm{C}_{d(s1)(22)}=\left(\bm{I}_{g}\otimes\bm{F}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$. Here, $\bm{A}^{*}=\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\bm{Z_{1}}\right)\bm{\mathit{\Sigma}}^{-1/2}$. ∎
@@
$\bm{A}^{*}$ $=\bm{\mathit{\Gamma}}^{-1}\otimes\left(\bm{H}_{n}\otimes\bm{V}^{*}\right),$ (S.6)
where $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$, and $\bm{\mathit{\Gamma}}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal matrix.
@@
$\bm{\mathit{\Sigma}}$ $=\bm{\mathit{\Gamma}}\otimes\left(\bm{I}_{n}\otimes\bm{V}\right),$ (S.7)
where $\bm{V}$ is a known positive definite and symmetric matrix. $\bm{\mathit{\Gamma}}$ is a $g\times g$ positive definite and symmetric matrix with non-zero off-diagonal elements.
@@
$\begin{split}\bm{\mathit{\Sigma}}^{-1}&=\bm{\mathit{\Gamma}}^{-1}\otimes\left(\bm{I}_{n}\otimes\bm{V}^{-1}\right),\\ \bm{\mathit{\Sigma}}^{-1}\bm{Z_{1}}&=\bm{\mathit{\Gamma}}^{-1}\otimes\left(\bm{I}_{n}\otimes\bm{V}^{-1}\right)\bm{X_{1}},\text{ and}\\ \bm{Z}^{{}^{\prime}}_{\mathbf{1}}\bm{\mathit{\Sigma}}^{-1}\bm{Z_{1}}&=\bm{\mathit{\Gamma}}^{-1}\otimes\bm{X}^{{}^{\prime}}_{\mathbf{1}}\left(\bm{I}_{n}\otimes\bm{V}^{-1}\right)\bm{X_{1}}.\end{split}$ (S.8)
From Bose and Dey, (2009), we get that $\left[\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)pr^{\perp}\left(\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\bm{X_{1}}\right)\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\right]=\bm{H}_{n}\otimes\bm{V}^{*}$, where $\bm{V}^{*}=\bm{V}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}^{-1}\bm{J}_{p\times p}\bm{V}^{-1}$. Thus from (S.8) and using the expression $\bm{A}^{*}=\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\bm{Z_{1}}\right)\bm{\mathit{\Sigma}}^{-1/2}$ as given in Lemma S.1, we get that for $g>1$
$\bm{A}^{*}$ $=\bm{\mathit{\Gamma}}^{-1}\otimes\left[\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)pr^{\perp}\left(\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\bm{X_{1}}\right)\left(\bm{I}_{n}\otimes\bm{V}^{-1/2}\right)\right]$
$=\bm{\mathit{\Gamma}}^{-1}\otimes\left(\bm{H}_{n}\otimes\bm{V}^{*}\right).$
@@
$=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{Z_{1}}&\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$ (S.10)
@@
$\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$ and $\bm{C}_{{d(s2)}(22)(1)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\bm{H}_{t}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$. Here, $\bm{A}^{*}=\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\bm{Z_{1}}\right)\bm{\mathit{\Sigma}}^{-1/2}$.
@@
$\bm{C}_{d(s2)}$ $=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\begin{bmatrix}\bm{Z_{1}}&\bm{I}_{2}\otimes\bm{F}_{d}\end{bmatrix}\right)\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right).$ (S.12)
From (S.12) and using the result $pr^{\perp}\left(\begin{bmatrix}\bm{\mathit{\zeta}_{1}}&\bm{\mathit{\zeta}_{2}}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)-pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)\bm{\mathit{\zeta}_{2}}\left(\bm{\mathit{\zeta}}^{{}^{\prime}}_{\mathbf{2}}pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)\bm{\mathit{\zeta}_{2}}\right)^{-}\times$
$\bm{\mathit{\zeta}}^{{}^{\prime}}_{\mathbf{2}}pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)$ (Kunert, 1983), we get
@@
where $\bm{C}_{d(s2)(11)}=\left(\bm{I}_{2}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$, $\bm{C}_{d(s2)(12)}=\bm{C}^{{}^{\prime}}_{d(s2)(21)}=\left(\bm{I}_{2}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\right)$ and $\bm{C}_{d(s2)(22)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\right)$. Here, $\bm{A}^{*}=\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\bm{Z_{1}}\right)\bm{\mathit{\Sigma}}^{-1/2}$.
Since $\bm{F}_{d}\bm{{1}}_{t}=\bm{{1}}_{n}\otimes\begin{bmatrix}0&\bm{{1}}_{p-1}^{{}^{\prime}}\end{bmatrix}^{{}^{\prime}}=\bm{P}\begin{bmatrix}0&\bm{{1}}_{p-1}^{{}^{\prime}}\end{bmatrix}^{{}^{\prime}}$, we get $\left(\bm{I}_{2}\otimes\frac{1}{t}\bm{F}_{d}\bm{J}_{t\times t}\right)=\left(\bm{I}_{2}\otimes\bm{P}\right)\times$
$\left(\bm{I}_{2}\otimes\frac{1}{t}\begin{bmatrix}0&\bm{{1}}_{p-1}^{{}^{\prime}}\end{bmatrix}^{{}^{\prime}}\bm{{1}}_{t}^{{}^{\prime}}\right)$. Hence, $\bm{I}_{2}\otimes\frac{1}{t}\bm{F}_{d}\bm{J}_{t\times t}$ belongs to the column space of $\bm{I}_{2}\otimes\bm{P}$. So
$\bm{\mathit{\Sigma}}^{-1/2}\left(\bm{I}_{2}\otimes\frac{1}{t}\bm{F}_{d}\bm{J}_{t\times t}\right)$ belongs to the column space of $\bm{\mathit{\Sigma}}^{-1/2}\bm{Z_{1}}$. Thus from (S.12), we can prove (S.10).
Further using the result $pr^{\perp}\left(\begin{bmatrix}\bm{\mathit{\zeta}_{1}}&\bm{\mathit{\zeta}_{2}}\end{bmatrix}\right)=pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)-pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)\bm{\mathit{\zeta}_{2}}\left(\bm{\mathit{\zeta}}^{{}^{\prime}}_{\mathbf{2}}pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)\bm{\mathit{\zeta}_{2}}\right)^{-}\bm{\mathit{\zeta}}^{{}^{\prime}}_{\mathbf{2}}pr^{\perp}\left(\bm{\mathit{\zeta}_{1}}\right)$ (Kunert, 1983), we get
@@
where $\bm{C}_{{d(s2)}(11)(1)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{T}_{d}\right)$, $\bm{C}_{{d(s2)}(12)(1)}=\bm{C}^{{}^{\prime}}_{{d(s2)}(21)(1)}=\left(\bm{I}_{2}\otimes\bm{T}_{d}^{{}^{\prime}}\right)\bm{A}^{*}\times$
$\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$ and $\bm{C}_{{d(s2)}(22)(1)}=\left(\bm{I}_{2}\otimes\bm{F}^{{}^{\prime}}_{d}\bm{H}_{t}\right)\bm{A}^{*}\left(\bm{I}_{2}\otimes\bm{F}_{d}\bm{H}_{t}\right)$. ∎
@@
$\bm{A}^{*}$ $=\begin{bmatrix}\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{1}&-\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\\ -\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}&\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{4}\end{bmatrix},$ (S.15)